OpenStudy (anonymous):

dy/dx = x^2(1-y) and f(0)=2, find the particular solution

OpenStudy (idku):

A similar example: \(\displaystyle dy/dx = 5x^4(2-y) ,~~~~~~~~~f(1)=1 \) Solution to this example: Separating variables and integrating: \(\displaystyle dy/dx = 5x^4(2-y) \) \(\displaystyle [1/(2-y)]~ dy = 5x^4~dx \) \(\displaystyle \int [1/(2-y)]~ dy = \int 5x^4~dx \) \(\displaystyle \ln|2-y| = x^5+C \) Determining C, by plugging initial value(s): \(\displaystyle \ln|2-1| = 1^5+C \) \(\displaystyle \ln(1) =1+ C \) \(\displaystyle C=-1\) Plugging the determined C into the solution: \(\displaystyle \ln|2-y| = x^5-1 \)

OpenStudy (anonymous):

wouldn't the integral of 1/(2-y) result in a negative?

OpenStudy (anonymous):

@idku

OpenStudy (bobo-i-bo):

yes you are right @gabyr5. @idku made a mistake and it should be \[-ln|2-y|\]